OpenStudy (anonymous):

Which of the following is not an element of the following set: A: {positive integers divisible by 3} 3 21 14 3

OpenStudy (swissgirl):

which one cant be divided by 3

OpenStudy (mathmate):

Integers are divisible when the sum of the digits is divisible by 3, and are not divisible when the sum is not.

OpenStudy (anonymous):

14?

OpenStudy (swissgirl):

3/3=1 21/3=7 14/3=? 3/3=1 yes

OpenStudy (anonymous):

thanks (:

OpenStudy (swissgirl):

3/3=1 21/3=7 14/3=? 3/3=1